OpenStudy (anonymous):

What are the exact solutions of x2 + 5x + 2 = 0? x = the quantity of negative 5 plus or minus the square root of 33 all over 2 x = the quantity of 5 plus or minus the square root of 33 all over 2 x = the quantity of negative 5 plus or minus the square root of 17 all over 2 x = the quantity of 5 plus or minus the square root of 17 all over 2

OpenStudy (mathmate):

use the quadratic formula. For \( x^2+5x+2=0\), a=1 b=5 c=2 where a,b and c are coefficients of the decreasing powers. The solution to the above equation is then \(\huge x = \frac{-b\pm\sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

I got D am I right?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

It's close but not quite the right one. If you notice that x=-b plus or minus... and that b=+5, then D is not the right answer.